Solve by the method of variation of parameters y"-y=xe^x/2
1
Expert's answer
2022-01-06T06:47:57-0500

Solution:

"y^{\prime \prime}-y=\underbrace{x e^{x/ 2}}_{g(x)} \text {. }"

The auxiliary equation is given by:

"\begin{aligned}

& D^{2}-1=0 \\

&\Rightarrow D^{2}=1 \Rightarrow \quad D=\pm 1

\end{aligned}"

complementary equation:"\quad y_{c}=c_{1} \underbrace{e^{ x}}_{y_{1}}+c_{2} \underbrace{e^{- x}}_{y_{2}}"

By variation of parameters,

Wronskian of the 2 functions is:


"W=\left|\begin{array}{ll}

y_{1} & y_{2} \\

y_{1}^{\prime} & y_{2}^{\prime}

\end{array}\right|=\left|\begin{array}{ll}

e^{x} & e^{-{x}} \\

 e^{x} & - e^{- x}

\end{array}\right|"

"=-e^0-e^0
\\=-1-1
\\=-2"

Now, the particular solution is given by:

"\begin{aligned}

y_{p}(x) &=y_{1} \int \frac{y_{2} g(x)}{W} d x+y_{2} \int \frac{y_{1} g(x)}{W} d x . \\

&=e^{{x}} \int \frac{e^{-{x}} \cdot x e^{{x/2}} d x}{-2}+e^{-{x}} \int \frac{e^{{x}} \cdot x e^{x / 2}}{-2} d x \\



\end{aligned}"

"=-\dfrac12 \int xe^{x / 2} d x-\dfrac12 \int xe^{x / 2} d x \\
=-\dfrac12 \int xe^{x / 2} d x\\"

On integrating by parts,

"y_p(x)=-\dfrac12[4\left(\frac{1}{2}e^{\frac{x}{2}}x-e^{\frac{x}{2}}\right)]
\\=-2\left(e^{\frac{x}{2}}x-e^{\frac{x}{2}}\right)
\\=2e^\frac{x}{2}-2xe^\frac{x}{2}"

"\therefore" The general solution is given by:

"y=y_{c}+y_{p}
\\=c_{1} e^{{x}}+c_{2} e^{-{x}}+2e^\frac{x}{2}-2xe^\frac{x}{2}"